A sperm cell of Equus caballus (horse) contains 32 chromosomes in its nucleus. a. How many DNA molecules would you expect in the daughter cells that result from a mitotic division of an epithelial cell of the same organism? b. How many DNA molecules does an epithelial cell of a horse contain in each one of the following phases? i. Start of interphase ii. End of interphase iii. Start of M phase iv. Metaphase c. What is the total weight of the genome of a horse epithelial cell culture that consists of 106 cells that are all in the metaphase, given that the genome of one haploid cell weighs 35ngr? d. If the length of the genome of a horse epithelial cell during metaphase is 4 x 109 base pairs, determine the length of the genome in a haploid cell of the same organism. e. How many pairs of homologous chromosomes are there in the nucleus of a somatic cell that was isolated from a male horse?
Answer #1

A sperm cell of Equus caballus (horse) contains 32 chromosomes in its nucleus.

a. How many DNA molecules would you expect in the daughter cells that result from a mitotic division of an epithelial cell of the same organism?

Answer: If the sperm cell has 32, then the somatic cell [epithelial cell] would have 2N, i.e. = 2 (32) = 64 chromosomes.

b. How many DNA molecules does an epithelial cell of a horse contain in each one of the following phases?

i. Start of interphase = 2N = 2 (32) = 64

ii. End of interphase = 4N = 4 (32) = 128

iii. Start of M phase = 4N = 4 (32) = 128

iv. Metaphase = 4N = 4 (32) = 128

c. What is the total weight of the genome of a horse epithelial cell culture that consists of 106 cells that are all in the metaphase, given that the genome of one haploid cell weighs 35ngr?

Answer: Mass of 1 haploid cell = 35 nano grams

Mass of 1 epithelial cell = 2 (35) = 70 nano grams

Mass of 106 cells = 70 (106) = 7 * 107 nano grams

d. If the length of the genome of a horse epithelial cell during metaphase is 4 x 109 base pairs, determine the length of the genome in a haploid cell of the same organism.

Answer: During the metaphase the size of the genome of a cell becomes 4 times the amount of genome in a haploid cell.

Length of genome in haploid cell = 4 x 109 / 4 =109.

e. How many pairs of homologous chromosomes are there in the nucleus of a somatic cell that was isolated from a male horse?

Answer: In a somatic cell, there are two pairs of homologous chromosomes.